# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ I.3 A full-stack hardware–software architecture for high-performance quantum computation
In order to tackle some of the key technical challenges listed above, here we introduce a heterogeneous quantum–classical full-stack hardware and software system architecture Humble et al. (2021). We outline how one could adopt existing semiconductor ecosystems and conventional high-performance infrastructures to build such an architecture, which is schematically illustrated in Figure 1 (see also Section V). At the highest layer, an HPC programming environment is extended to include a quantum accelerator API consisting of an interface library for seamless invocation of quantum kernels, an adaptive circuit knitting hypervisor for efficient quantum workload partitioning and distribution, and a hybrid quantum–classical compiler. A hybrid quantum–classical workload manager ensures optimal quantum resource utilization in a multi-user environment. To support fault-tolerant quantum computation, a compiler, emulator, assembler, and a real-time decoder together use known hardware noise profiles to synthesize optimized fault-tolerant circuits to solve the problem at hand. Calibration and control of quantum resources use specialized hardware that are integrated with the HPC. Heterogeneous coprocessors—including CPU/GPUs, quantum processing units (QPU), probabilistic processing units (PPU)—allow the system to partition a particular problem into subproblems that can be sent to the appropriate coprocessors.
@@
    Control hardware (Section II.3) can be realized and benchmarked by existing room-temperature control electronics for up to a thousand qubits. The 20k-qubit scale could be achieved by a combination of i) wafer-scale integration, ii) high-density cables and interconnects, iii) a moderate level of time and frequency division multiplexing (1:4 or 1:8), and iv) low-power, high-density digital-to-analog front-end development. Finally, to reach the 1M-qubit scale, dedicated and integrated CMOS could be developed that may operate in cryogenic temperatures. In addition, tight integration between the control hardware and the compute resources of the HPC system must be developed to allow for efficient calibration workflows that optimize and stabilize fidelities while not limiting uptimes and system utilization.
@@
    For fault-tolerant quantum computing (Section III) an error-correction decoder (Section III.6) can be tested using an HPC-accelerated FTQC emulator (Sections III.2 and III.5 and Appendix E) to synthesize syndrome measurements of QEC codes. This test data could then be directed into the control hardware and used to benchmark the decoding CPU/GPU hardware and software in real-time emulation. For utility-scale FTQC, the decoding software system may require incorporating all ideas of distributed, hierarchical, and moving-window decoding to support 1M+ qubits.
@@
    The quantum computer can be integrated with conventional HPC infrastructures in six different layers (Section V): heterogeneous quantum–classical coprocessors, adaptive circuit knitting, FTQC compilers, distributed decoders, calibration, and control. A dedicated ultra-low latency, real-time QEC network that includes the control hardware, the decoder hardware, and the logical circuit orchestration hardware could be utilized to execute the FTQC workflow.
@@ II.1 Qubit fabrication
Researchers often report the performance of their best superconducting qubits, as opposed to more meaningful metrics such as an average, or as more appropriate for systems engineering, the worst device performance. A current challenge is thus to fairly compare approaches and build reliable models for improving coherence. In Table 1, we derived the “Target” hardware parameters needed to achieve an FTQC error suppression rate for practical quantum advantage. In particular, we introduce a new “tailedness” metric that describes the distribution in qubit quality. For more information on how these metrics where obtained, refer to Appendix B.
@@
Table 1: Hardware specifications for three sets of parameters: baseline, target, and desired hardware. The baseline set represents the state-of-the-art values; the target set is envisioned to be a promising near-term goal; the desired set of synthetically generated hardware specifications corresponds to a noise model with about twice the error suppression rate, $\Lambda$, of the target hardware extracted from the exponential suppression law $\mu d^{2}\Lambda^{-(d+1)/2}$ for quantum memory. Our benchmarking studies resulted in $\Lambda\approx 2.34$ for the baseline set and $\Lambda\approx 9.3$ for the target set, respectively. The desired model is specified by the value $\Lambda\approx 18$ and is therefore also referred to as the “$\Lambda_{18}$ model” in this paper. The $T_{1}$ tailedness characterizes the weight of poor-quality qubits with respect to variations in coherence times across the qubit chip. As discussed in Section III.4, the standard deviation is used as the metric for tailedness in this paper, while the effects of higher moments (such as skewness and kurtosis) can also be crucial given that realistic distributions of $T_{1}$ values have significantly heavier tails compared with the associated approximating Gaussian distributions.
@@
Based on this challenge, a systems engineering principle that can be used to guide qubit metrology is “the worst 1% of devices determine system performance.” For quantum computers, this is a useful design rule since an $\sim$1qubit dropout rate for the surface code will cause the system to degrade or fail. Our goal is thus similar to fabricating complex CMOS electronics: make every qubit identically good. Our plan is based on the categories of quality, quantity, speed, and connectivity introduced in the previous section, but organized according to hardware subsystems. Fortunately, we have found that many of the system engineering constraints can be solved concurrently. For example, we will explain how qubits can be fabricated in a manner to simultaneously improve both quality and scaling.
@@
Fabrication. We do not assume this academic process is good enough, it only shows that our models and fabrication plans for improving qubits are on track. An example is our insights on TLS defects, a significant issue for present-day superconducting qubits. Two-level systems introduce sparse defects at random frequencies which lowers coherence for a significant fraction (1–10%) of qubits, thus they cannot be statistically avoided in large systems. Preliminary data of our fabrication process shows a much lower density of these TLS defects, and even a significant tightening of the spread of $T_{1}$ coherence times.
@@ II.2 Wafer-scale integration
Measurement and readout. Another scaling bottleneck is qubit measurement and readout, as typical systems require circulators and parametric amplifiers that have a volume of $\sim$1 cm3 or more. Qolab plans to use a readout technology that can be readily integrated into the qubit or wiring wafer, as described in Figs. 7 and 8. Readout using the Josephson photomultiplier, developed by McDermott Opremcak et al. (2020), has achieved acceptable fidelity of 99% and can further be improved. Because the measurement forms a classical state on the integrated detector, it can be readout in a multiplexed manner with a simple superconducting SLUG (SQUID) amplifier at 3 K, without needing circulators. CMOS drivers at 3 K have been developed to bias the large number of SLUG amplifiers.
@@
Scaling through tiling. For quantum computers larger than 20k qubits, one could create baseline system that uses subsytem tiling, as illustrated in Figure 9. Here, each qubit wafer is precisely mounted onto an invar frame so the modules can be mechanically assembled with precise capacitive coupling between the wafers. These “edge couplers” communicate between each tile and can have higher error rates than qubits within each tile.
Figure 9: Tiling of multiple qubit subsystems. Each subsystem has 20 k qubits, as described in Figure 5.
Figure 9 illustrates linear tiling, but a serpentine pattern is possible for a more compact area. A system with 5–10 tiles can fit into a large dilution refrigerator using present-day designs. Scaling up to many more tiles would clearly require new designs for a large cryostat. Optical connects Mirhosseini et al. (2023) could provide an alternative solution allowing easier scaling with modular cryostats. In this case, it would be particularly useful to separate the $T$-gate distillation factory from the main processor. These ideas will be incorporated as soon as optical to microwave quantum transducers are available. Resource estimates for both capacitive interconnects and optical interconnects is described in Section III.7
@@ II.3 Control hardware
Generating pulses within the coherence time of the qubits, acquiring qubit measurements, and processing these measurements for real-time feedback and efficient data transfer requires a unique digital processor architecture, which we refer to as the pulse processor unit. Then, the system’s analog front-end, which includes the digital-to-analog and analog-to-digital converters as well as the analog signal chains (amplifiers, filters, attenuators, etc.), generates and acquires analog signals. As the system scales, maintaining analog performance, particularly in terms of noise, stability, and cross-talk, becomes crucial for achieving the fidelity targets necessary for NISQ computing or QEC.
Scaling quantum control. With current qubit quality and scale up to 1k qubits, the control system architecture should focus on performance and flexibility. Such performance and flexibility is needed so the control system does not limit overall performance or the speed of testing, research, and development iterations, even at the cost of overall design. As the system scales, the density of control electronics and wiring presents challenges in terms of space and heat dissipation. Moreover, cost per qubit control becomes a significant issue. Hence, once the desired fidelities are achieved and the requirements for errors and scale-up are well understood, some control system requirements can be relaxed. This would allow optimizing for cost, power consumption, and size while maintaining target fidelities. To reach 20k qubits, one could employ a combination of low-power, high-density, room temperature analog front-end development in conjunction with cryo-CMOS components for moderate 1:4, or 1:8 frequency and/or time division multiplexing. To reach 1M qubits, one could develop dedicated, integrated CMOS that may operate in cryogenic temperatures. Effective co-design of room-temperature and cryogenic control electronics will be critical for seamless integration as the system scales.
Another possibility for scalable qubit control might arise from our advanced fabrication, assuming an improvement in qubit quality and reproducibility: if the qubits can be fabricated nearly identically and with a low probability of TLS dropouts, then a single control signal can be split to many qubits, with simple variable amplitude and phase adjusters at 3 K to fine-tune signals. This will enable an integrated and low dissipation solution to control.
@@
Room temperature and cryogenic control. As quantum systems scale, the benefits of cryogenic control increases. Wafer-scale integration enables the connection of flex cables at 3 K, where a maximum heat load of 41 W is expected to be sufficient for flex cables connecting control at 77 K or 300 K. We identify several trade-offs between room-temperature and cryogenic control at 77 K or 3 K:
@@
    Power consumption: Operating at cryogenic temperatures with lower $V_{dd}$ and reduced signal amplitude, proportional to the reduced thermal noise, leads to significantly lower power consumption—potentially orders of magnitude lower compared to room-temperature systems.
@@
    Size and system complexity: Cryogenic control, such as at 77 K, results in a more compact system, housed entirely within the cryogenic refrigerator. This not only eliminates the need for multiple racks but also reduces the number of flex cables required between the cryogenic and room-temperature stages, leading to a more efficient setup.
@@
    Process and cost: Room-temperature electronics utilize established CMOS processes, benefiting from mature manufacturing ecosystems, stability, and relatively low costs. Cryogenic control, however, requires more specialized processes like fully depleted silicon-on-insulator (FDSOI), optimized for low power. Broader adoption of these processes in industries like aerospace could drive down costs and improve availability, facilitating wider adoption.
@@
Based on the current analysis, low-power digital-to-analog converters (DAC) along with their analog chains are crucial components for enabling low-temperature control. Achieving a power target below 1 mW per channel could enable overall analog and digital power consumption of approximately 100 W at 77 K, sufficient for controlling a 20k qubit system. Once the target qubit quality is achieved, a comprehensive end-to-end trade-off analysis should be conducted to determine the optimal architecture for practical implementations, considering both cryogenic and room-temperature control.
@@
For systems of 10k–100k qubits, local optimization of the QPU, packaging, room-temperature control, and cryogenic electronics become inadequate. Instead, the control system must be optimized as part of a holistic solution. For instance, channel crosstalk and pre-distortion issues can be addressed either at the QPU and packaging level or within the control system. A localized optimization would yield suboptimal results, whereas a co-design approach can produce an optimal solution. Control filters should be tailored to the package and QPU channels, while crosstalk can be mitigated at the CPU and package level, ensuring the control architecture is aligned to correct for the residual errors. Similarly, analog characteristics of the control system such as phase stability should be optimized based on gate implementation to meet the system’s architectural requirements and prevent unnecessary complexity and cost.
@@
In large-scale systems, both control systems and quantum processors are prone to failures. Consequently, it is essential to incorporate mechanisms for testing functionality at both the system and component levels. Decoder results from benchmark circuits can offer insight into the overall operational status of the system. Additionally, benchmarks focusing on one or few surfaces provide a means to diagnose localized failures. Benchmarks evaluating the fidelity of physical qubits can offer insights into qubit calibration status as well as potential hardware issues. Furthermore, component-specific benchmarks such as those for amplifiers are necessary to isolate and identify exact points of failure. These benchmarks should be optimized for execution in minimal time to enable frequent testing, and may also be used to trigger calibration procedures when necessary. Because large-scale control hardware is susceptible to local errors, robust management and monitoring features—encompassing telemetry and self-testing—are required to detect, diagnose, and address such issues effectively. Upon detection of an error, the logical circuit orchestrator is notified, ensuring the platform continues to function while debugging and calibration processes are applied to resolve the faults.
@@
To facilitate debugging and hardware updates while maintaining an operational system, the control platform is structured with multiple clusters. Each cluster is managed independently while remaining synchronized with the others, thus enabling the isolation of individual clusters for maintenance or error handling without disrupting the broader system.
@@
Figure 12: Illustration of an FTQC platform. Multiple quantum control clusters are connected to multiple QPUs. The platform includes optional cryogenic control located adjacent to the QPU. The quantum low-latency network connects the clusters and acceleration servers that run the QEC decoders, logical circuit orchestration and calibration and optimal control routines. A high-performance user network connects the control and servers and is used for tight integration with HPC compute resources.
Control system integration with classical compute resources. By their nature, quantum computing workloads require quantum–classical iterations. This includes calibration workflows, NISQ hybrid applications, and quantum error correction. Quantum–classical iterations require a low-latency, high-throughput interface with the right division of responsibilities that allows transferring a compact representation of the data across systems. For applications that require significant compute resources, tight integration to HPC clusters is required to efficiently utilize the quantum and classical resources.
Calibration impacts qubit fidelity and plays a critical role in determining the error correction code distance, which in turn directly affects the scalability and complexity of the decoding process. A significant challenge is variability in qubit fidelity; the system’s overall performance is often constrained by the lowest-performing outlier qubits. To maximize the performance of these outliers, advanced calibration strategies such as optimal control, reinforcement learning, demonstrated in Figure 11, and model-based simulations may be used. To meet the fidelity goals for large quantum processors, it will be necessary to have scalable and efficient calibration routines that enable concurrent calibration across qubits. Reducing the calibration time allows for repeating the calibration more frequently and allocating a larger portion of the time to optimizing outlier qubit performance. Figure 11 demonstrates the fidelity improvement from frequent calibration, tracking the drift in a qubit frequency. A key requirement for calibration flows is minimal execution overhead and feedback latency. A typical calibration node requires executing thousands or more shots. A shot is structured with initialization, executing a pulse sequence, and measurement, and typically takes hundreds of nanoseconds. As such, the calibration routine overhead should be minimized accordingly to few milliseconds from loading to gathering the measurements statistics.
Architecture of a quantum–classical integrated platform for FTQC. FTQC workflows require a quantum–classical integrated platform to execute the logical circuit, orchestrate QEC decoding and surface operations, and control physical qubits. The main components of the platform illustrated in Figure 12 include the quantum control, QEC acceleration servers, and dilution fridges that hold the quantum processor and cryogenic control. The quantum control is divided into clusters; typically each cluster controls a set of surfaces, and the independent clusters simplify the control and provide resiliency to errors while maintaining synchronized execution. The QEC acceleration servers are responsible for real-time decoding and logical circuit orchestration. The servers leverage general-purpose GPU/CPU systems with the option for additional ASIC or FPGA accelerators. The low latency QEC network provides an efficient interface for transferring syndromes and surface operations. A high-performance user network is used for data and program loading as well as connectivity to HPC resources required, for example, for logical circuit compilation.

### v2

@@ I.3 Exponential scaling progress: A mirage or reality?
Figure 2: Architecture diagram of a quantum–classical full-stack solution. Extensions within the HPC programming environment include a quantum interface library for seamless invocation of quantum kernels, an adaptive circuit knitting hypervisor for efficient quantum workload partitioning and distribution, and quantum compiler and runtime extension for performant quantum circuit compilation. A customized hybrid workload manager ensures maximal quantum resource utilization in a multi-user environment. For fault-tolerant quantum computation, a compiler, emulator, assembler, and real-time decoder work together to use hardware noise profiles to synthesize optimized fault-tolerant circuits. Calibration and control of quantum resources use specialized hardware and are integrated with HPC. At the hardware layer, heterogeneous coprocessors include CPUs/GPUs, quantum processing units (QPU), probabilistic processing units (PPU), FPGA and custom-design ASICs with high-speed low-latency scale-up interconnect.
As we have outlined in this position paper, to maintain the rate of exponential growth it is important to identify all major technical challenges that are impeding progress and devise mitigation strategies. Such strategies include taking a radically different approach to qubit fabrication and developing a full-stack system integration with heterogeneous high-performance computing infrastructures.
I.4 A full-stack hardware–software architecture for high-performance quantum computation
Addressing key technical challenges listed above requires a heterogeneous full–stack quantum–classical hardware and software system architecture Svore et al. (2006); Humble et al. (2021). To this end, we outline how one could adopt existing semiconductor ecosystems and conventional high-performance infrastructures to build such an architecture, which is schematically illustrated in Figure 2 (see also Section V). At the highest layer, an HPC programming environment is extended to include a quantum accelerator API consisting of an interface library for seamless invocation of quantum kernels, an adaptive circuit knitting hypervisor for efficient quantum workload partitioning and distribution, and a hybrid quantum–classical compiler. A hybrid quantum–classical workload manager ensures optimal quantum resource utilization in a multi-user environment. To support fault-tolerant quantum computation, a compiler, emulator, assembler, and a real-time decoder together use known hardware noise profiles to synthesize optimized fault-tolerant circuits to solve the problem at hand. Calibration and control of quantum resources use specialized hardware that are integrated with the HPC. Heterogeneous coprocessors—including CPU/GPUs, quantum processing units (QPU), probabilistic processing units (PPU)—allow the system to partition a particular problem into subproblems that can be sent to the appropriate coprocessors.
Figure 3 shows a schematic architecture layout for future accelerated quantum supercomputers. Here, the key approach is a tight integration of QPUs as specialized accelerators within AI supercomputers. This approach combines the capabilities of quantum and classical computation, with AI superchips executing computationally demanding classical tasks and QPUs handling specialized quantum tasks. AI supercomputing can also play a complementary role in enhancing the capabilities of QPUs by executing, for example, QEC, state preparation, and optimization tasks. A low-latency interconnect between AI superchips and QPUs is a critical component of HPC–quantum integration in executing QEC. In addition, classical interconnects between AI superchips allow scalable parallel processing, while quantum interconnects between QPUs facilitate entanglement distribution and resource sharing for distributed quantum computing. By tightly coupling classical and quantum components through optimized interconnects, this architecture provides a scalable foundation for future quantum supercomputers.
@@ I.4 A full-stack hardware–software architecture for high-performance quantum computation
Figure 3: Architecture of two nodes in an accelerated quantum supercomputer. AI superchips are coupled to QPUs via low-latency quantum–classical interconnects and quantum control hardware. AI superchips communicate via classical HPC interconnects, while QPUs are connected by quantum interconnects.
@@
    Control hardware (Section II.3) can be realized and benchmarked by existing room-temperature control electronics for up to a thousand qubits. The 20k-qubit scale could be achieved by a combination of i) wafer-scale integration, ii) high-density cables and interconnects, iii) a moderate level of time and frequency division multiplexing (1:4 or 1:8), and iv) low-power, high-density digital-to-analog front-end development. To reach the 1M-qubit scale, dedicated and integrated CMOS may need to be developed to operate at cryogenic temperatures. In addition, tight integration between control hardware and compute resources of the HPC system must be developed to allow for efficient calibration workflows that optimize and stabilize fidelities while not limiting uptimes and system utilization.
@@
    For fault-tolerant quantum computing (Sections III and IV and Appendices A, B, C, D and E) an error-correction decoder (Section III.6) can be tested using an HPC-accelerated FTQC emulator (Sections III.2 and III.5 and Appendix E) to synthesize syndrome measurements of QEC codes. This test data could then be directed into the control hardware and used to benchmark the decoding CPU/GPU hardware and software in real-time emulation. For utility-scale FTQC, the decoding software system may require incorporating all ideas of distributed, hierarchical, and moving-window decoding to support 1M+ qubits.
@@
    The quantum computer can be integrated with conventional HPC infrastructures in six different layers (Section V and Appendices F and G): heterogeneous quantum–classical coprocessors, adaptive circuit knitting, FTQC compilers, distributed decoders, calibration, and control. A dedicated ultra-low latency, real-time QEC network that includes the control hardware, the decoder hardware, and the logic circuit orchestration hardware could be utilized to execute the FTQC workflow.
@@ II.1 Qubit fabrication
Researchers often report the performance of their best superconducting qubits, as opposed to more meaningful metrics such as an average, or as more appropriate for systems engineering, the worst device performance. A current challenge is thus to fairly compare approaches and build reliable models for improving coherence. In Table 1, we derived the target hardware parameters needed to achieve an FTQC error suppression rate for practical quantum advantage. In particular, we introduce a new tailedness metric that describes the distribution in qubit quality. For more information on how these metrics were obtained, refer to Appendix B.
@@
Table 1: Hardware specifications for three sets of parameters: baseline, target, and desired hardware. The baseline set represents the state-of-the-art values; the target set is envisioned to be a promising near-term goal; the desired set of synthetically generated hardware specifications corresponds to a noise model with about twice the error suppression rate, $\Lambda$, of the target hardware extracted from the exponential suppression law $\mu d^{2}\Lambda^{-(d+1)/2}$ for quantum memory. Our benchmarking studies resulted in $\Lambda\approx 2.34$ for the baseline set and $\Lambda\approx 9.3$ for the target set, respectively. For comparison, the exponential suppression rate $\bar{\Lambda}$ following the commonly adopted model $\bar{\mu}d\bar{\Lambda}^{-(d+1)/2}$ is also provided (see Section III.2 for more details). The $T_{1}$ tailedness characterizes the weight of poor-quality qubits with respect to variations in coherence times across the qubit chip. As discussed in Section III.4, the standard deviation is used as the metric for tailedness in this paper, while the effects of higher moments (such as skewness and kurtosis) can also be crucial given that realistic distributions of $T_{1}$ values have significantly heavier tails compared with the associated approximating Gaussian distributions.
@@
Based on this challenge, qubit metrology should be driven by system performance metrics based on the worst 1% of devices. For quantum computers, this is a useful design rule since an $\sim$1qubit dropout rate for the surface code will cause the system to degrade or fail. Our goal is thus similar to fabricating complex CMOS electronics: make every qubit identically good. Our plan is based on the categories of quality, quantity, speed, and connectivity introduced in the previous section, but organized according to hardware subsystems. Fortunately, we have found that many of the systems engineering constraints can be solved concurrently. For example, we will explain how qubits can be fabricated in a manner to simultaneously improve both quality and scaling.
@@
TLS modeling. Developing sophisticated machine-learned models to optimize and characterize the structural properties of the Si-Al interface, with particular attention placed on potential oxygen incorporation, will create the best possible conditions for superconducting -based qubits. Leveraging a moment tensor potential trained on Kohn-Sham DFT data, the model can elucidate how variations in inter-facial structure directly influence the electronic behavior of aluminum, a critical factor in superconducting applications. Additionally, NEGF methods can be employed to provide a thorough analysis of electronic properties, allowing for exploration of the impact of fabrication processes on device performance. By linking structural characteristics, process variables, and electronic outcomes, TLS modeling can drive innovation in the design of materials optimized for quantum computing environments.
Fabrication. Far from sufficient, the above academic manufacturing process shows that our models and fabrication plans for improving qubits are on track. This can be illustrated byour insights on TLS defects, a significant issue for present-day superconducting qubits. Two-level systems introduce sparse defects at random frequencies which lowers coherence for a significant fraction (1–10%) of qubits, thus they cannot be statistically avoided in large systems. Preliminary data of our fabrication process shows a much lower density of these TLS defects, and even a significant tightening of the spread of $T_{1}$ coherence times.
@@ II.2 Wafer-scale integration
Measurement and readout. Another scaling bottleneck is qubit measurement and readout, as typical systems require circulators and parametric amplifiers that have a volume of $\sim$1 cm3 or more. Qolab plans to use a readout technology that can be readily integrated into the qubit or wiring wafer, as described in Figs. 9 and 10. Readout using the Josephson photomultiplier, developed by McDermott Opremcak et al. (2020), has achieved acceptable fidelity of 99% and can further be improved. Because the measurement forms a classical state on the integrated detector, it can be readout in a multiplexed manner with a simple superconducting SLUG (SQUID) amplifier at 3 K, without needing circulators. CMOS drivers at 3 K have been developed to bias the large number of SLUG amplifiers.
@@
Scaling through tiling. For quantum computers larger than 20k qubits, one could create baseline system that uses subsystem tiling where, each qubit wafer is precisely mounted onto an invar frame so the modules can be mechanically assembled with precise capacitive coupling between the wafers. These “edge couplers” communicate between each tile and can have higher error rates than qubits within each tile.
In addition to linear tiling, serpentine pattern is possible for a more compact area. A system with 5–10 tiles can fit into a large dilution refrigerator using present-day designs. Scaling up to many more tiles would clearly require new designs for a large cryostat. Optical connects Mirhosseini et al. (2023) could provide an alternative solution allowing easier scaling with modular cryostats. In this case, it would be particularly useful to separate the $T$-gate distillation factory from the main processor. These ideas will be incorporated as soon as optical-to-microwave quantum transducers are available. Resource estimates for both capacitive interconnects and optical interconnects is described in Section III.7
@@ II.3 Control hardware
Generating pulses within the coherence time of the qubits, acquiring qubit measurements, and processing these measurements for real-time feedback and efficient data transfer require a unique digital processor architecture, which we refer to as the pulse processor unit. Then, the system’s analog front-end, which includes the digital-to-analog and analog-to-digital converters as well as the analog signal chains (amplifiers, filters, attenuators, etc.), generates and acquires analog signals. As the system scales, maintaining analog performance, particularly in terms of noise, stability, and crosstalk, becomes crucial to achieving the fidelity targets necessary for NISQ computing or QEC.
Scaling quantum control. With current qubit quality and scale up to 1k qubits, the control system architecture should focus on performance and flexibility. Such performance and flexibility are needed so the control system does not limit overall performance or the speed of testing, research, and development iterations, even at the cost of overall design. As the system scales, the density of control electronics and wiring presents challenges in terms of space and heat dissipation. Moreover, cost per qubit control becomes a significant issue. Hence, once the desired fidelities are achieved and the requirements for errors and scale-up are well understood, some control system requirements can be relaxed. This would allow optimizing for cost, power consumption, and size while maintaining target fidelities. To reach 20k qubits, one could employ a combination of low-power, high-density, room temperature analog front-end development in conjunction with cryo-CMOS components for moderate 1:4, or 1:8 frequency and/or time division multiplexing. To reach 1M qubits, one could develop dedicated, cryogenic integrated CMOS connected by a digital interface to the room-temperature control. Effective co-design of room-temperature and cryogenic control electronics will be critical for seamless integration as the system scales.
Another possibility for scalable qubit control might arise from our advanced fabrication, assuming an improvement in qubit quality and reproducibility: if the qubits can be fabricated nearly identically and with a low probability of TLS dropouts, then a single control signal can be split to many qubits, with simple variable amplitude and phase adjusters at 3 K to fine-tune signals.
@@
Room temperature and cryogenic control. As quantum systems scale, the benefits of cryogenic control increase. Wafer-scale integration enables the connection of flex cables at 3 K connecting control at 77 K or 300 K. We identify several trade-offs between room temperature and cryogenic control at 77 K or 3 K:
@@
    Power consumption: Operating at cryogenic temperatures with lower $V_{dd}$ and reduced signal amplitude, proportional to the reduced thermal noise, leads to significantly lower power consumption, potentially orders of magnitude lower compared to room temperature systems.
@@
    System complexity and size: Cryogenic control, such as at 77 K, results in a more compact system, housed entirely within the cryogenic refrigerator. This not only eliminates the need for multiple racks but also reduces the number of flex cables required between the cryogenic and room-temperature stages, improving overall efficiency.
@@
    Process and cost: Room-temperature electronics utilize established CMOS processes, benefiting from mature manufacturing ecosystems, stability, and relatively low costs. Cryogenic control, however, requires more specialized processes like fully depleted silicon-on-insulator (FDSOI), optimized for low power. Broader adoption of these processes in industries like aerospace could reduce costs and improve availability, facilitating wider adoption.
@@
Based on the current analysis, low-power digital-to-analog converters (DAC) along with their analog chains are crucial components for enabling low-temperature control. A power target of 1 mW per channel should allow the control to operate at 77 K. Once the target qubit quality is achieved, a comprehensive end-to-end trade-off analysis should be conducted to determine the optimal architecture for practical implementations, considering both cryogenic and room-temperature control.
@@
For systems of 10k100k qubits, local optimization of the QPU, packaging, room temperature control, and cryogenic electronics become inadequate. Instead, the control system must be optimized as part of a holistic solution. For instance, the predistortion capabilities of the control should be designed to address the given channels distortion and crosstalk should be addressed in the QPU and chip mount design with the control handling only the residual errors. Similarly, analog characteristics of the control system, such as phase stability, should be optimized based on gate implementation to meet the system’s architectural requirements and avoid unnecessary complexity and cost.
@@
Management and monitoring. In large-scale systems, both control systems and quantum processors are prone to failure. Consequently, it is essential to incorporate mechanisms for testing functionality at both the system and component levels. Decoder statistics can offer insight into the overall operational status of the system. Additionally, benchmarks on one or few surfaces provide a means to diagnose localized failures. Benchmarks evaluating the fidelity of physical qubits can offer information on qubit calibration status as well as potential hardware issues. Furthermore, component-specific benchmarks such as those for amplifiers are necessary to isolate and identify exact points of failure. These benchmarks should be optimized for execution in minimal time to enable frequent testing, and may also be used to trigger calibration procedures when necessary. Because large-scale control hardware is susceptible to local errors, robust management and monitoring features such as telemetry and self-testing are required to detect, diagnose, and address such issues effectively. Upon detection of an error, the logic circuit orchestrator is notified, ensuring continued platform function while debugging and calibration processes resolve the faults.
@@
To facilitate debug and control software updates while maintaining an operational system, the control platform is structured with multiple clusters. Each cluster is managed independently with the clusters operate in synchronization. This provides the isolation of individual clusters for maintenance or error handling without disrupting the broader system.
@@
Figure 13: Illustration of an FTQC platform. Multiple quantum control clusters are connected to multiple QPUs. The platform includes optional cryogenic control located adjacent to the QPU. The low-latency QEC network connects the clusters and acceleration servers that run the QEC decoders, logic circuit orchestration and calibration and optimal control routines. A high-performance user network connects the control and servers and is used for tight integration with HPC compute resources.